Determine the appropriate form of the particular solution for the following non-homogeneous linear differential equation with constant coefficients.J.(4) +9y = 5 + e (x – 3) + 4sin(3x). Ax + B + C sin(3x) + D cos(3x) + Exer Ax? + Bxe - 3x + Cxe3x + Det + Exet A + Bxe-3x
Answer #1

(4) y Given: +9y -5+ ex (1-3) + 4 sin(s) y (4) +94=5+ sex-30 + 4 Sin (3x) Auxillary equation is m4tama=o =) ma (m2 +9)=0 m